# Vulnerabilities of quantum key distribution systems in visible range

### v1

@@ Abstract
Spectral loopholes in the 1000-2100 nm range have been the focus of attention of quantum hacking research aimed at searching and elimination of vulnerabilities in quantum key distribution systems. In this work, we concentrate on shorter wavelengths ranged from 400 nm up to 800 nm and experimentally study spectra of insertion losses for a number of fiber optical components to assess potentials for their implementation as countermeasures. We show that efficiency of the elements commonly used as countermeasures against hacking attacks in the telecom range can be significantly impaired in the visible range. It is found that the vulnerabilities detected in this range play an increasingly important role for hacking strategies such as the induced-photorefraction attack whose efficiency improves when the wavelength decreases.
@@ I Introduction
The security of quantum key distribution (QKD) systems hinges upon the fundamental laws of physics. Over the past decades, significant progress has been made in theoretical security proofs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. However, many of these proofs rest on the assumption of perfect and trusted devices. Such trust assumption imposes a rather strong security requirement which is generally difficult to implement in practice. In real situations imperfect devices in QKD systems would introduce vulnerabilities that can be exploited by eavesdroppers to extract additional information about the transmitted bits during their attacks as it is considered in [11, 12, 13, 14, 15, 16]. Device-independent (DI) [17, 18, 19, 20, 21] and semi-device-independent [22, 23, 24, 25] protocols have emerged as a prominent area of research that provides a way to prevent side-channel attacks [26]. On the opposite side, another rapidly developing field which is often referred to as quantum hacking has recieved a considerable amount of attention [27].
@@
Emerging approaches in quantum hacking pose novel challenges for security analysis of QKD systems. Numerous vulnerabilities have been identified in prior works [28, 29, 30], underscoring the importance of enhancing security measures. Eavesdropper strategies in quantum hacking can broadly be classified under two basic types, active and passive, as well as combinations thereof. Active strategies involve the manipulation of optical parameters within the hardware, whereas passive strategies solely facilitate the acquisition of additional information.
@@
A quintessential illustration of a passive attack is the Trojan-horse attack which is also known as the large pulse attack [31, 32]. In this scenario, the eavesdropper (Eve) employs scanning pulses introduced into the systems of legitimate users (Alice and Bob) via the quantum channel, allowing for the extraction of distributed sequences from the reflected portions of the eavesdropper pulses back into the quantum channel. The key factors that determine the efficiency of this attack involve the transmittance of elements in two directions, the distinguishability of states along with countermeasures utilized in the system. An important point is that all these factors require the eavesdropper to define the optimal spectral range of wavelengths employed to perform attack.
@@
The well-known detector blinding attack (DBA) [33, 34, 35] exemplifies an active attack. The idea underlying the attack is that, under the action of high-power laser radiation (as compared to typical intensities used in QKD), the operating mode of single photon detectors (SPDs) based on avalanche photodiodes can be changed from the Geiger mode into the linear operating mode, so that the eavesdropper can manipulate the clicks registered by the SPDs. The efficiency of this DBA based quantum hacking approach crucially depends on the spectral sensitivity of the avalanche photodiode, the transmittance of optical elements connecting the quantum channel with the single photon avalanche diode (SPAD), and the countermeasures implemented into the QKD system. So, it turned out that, similar to the Trojan-horse attack, the spectral range is one of the factors that govern success of the DBA attack.
@@
Previous studies [36, 37, 38, 39] motivated by the THA, have mainly been focused on transmittance of fiber optical elements utilized in QKD systems at wavelengths ranged between 1000 nm and 2100 nm. For the 600-1000 nm range, the experimental results were reported in Refs. [39, 40].
@@
Note that, according to Ref. [46], successful implementation of the IPA requires low intensities starting from 3 nW. A comparison between these intensities and the maximal intensity permissible for insertion into a fiber without causing damage, which is typically around 9 W (this value is applicable to quartz single mode fibers without power damage level doping commonly used in fiber communications and may vary among manufacturers) leads to the conclusion that attenuation needed for passive defence against IPA for single pass through optical elements can roughly be estimated at about 100 dB. Thus, in order to assess the feasibility of the attack, it is necessary to measure and analyze insertion losses of individual optical elements on the side of Alice.
To this end, we study the insertion losses of different optical elements in the 400–800 nm spectral range, demonstrate the existence of potential vulnerabilities and suggest potential countermeasures. The paper is organized as follows.
@@
In Sec. II we describe our experimental technique employed to measure spectra of insertion losses of optical elements at wavelengths ranged from 400 nm to 800 nm. The experimental data for insertion losses measured for a variety of optical attenuators, wave-division multiplexing (WDM) and isolating components are presented in Sec. III. Finally, in Sec. IV, we discuss and summarize our results.
II Experimental setup and method
@@
Figure 1 shows the optical scheme of our experimental setup used to measure the insertion losses of optical elements in the visible range. In this scheme, a broadband plasma light source (XWS65, ISTEQ) indicated as the broadband light source (BLS) is utilized to produce light with wavelengths ranged from 250 nm up to 2500 nm. Referring to Fig. 1, the radiation from this source is directed into a single-mode fiber (OF) passing through a set of lenses (L), filters (F) and the objective (O) which is positioned on a three-axis stage to meet coupling and spectrum requirements. After the fiber, the focused light propagates through either a fiber or a fiber with the optical elements under test and is measured by the spectrometer (USB4000-UV-VIS-ES, Ocean Optics) with the 1.5 nm resolution.
@@
Figure 2: Insertion losses measured in the 400-800 nm range for (a) electro-mechanical and (b) electro-optical variable optical attenuators (blue and red solid lines indicate the zero-voltage curve and the case where the applied voltage is 5 V, respectively); (d) absorption based fixed attenuators (the curves for the first 20 dB FOA, second 20 dB FOA and 10 dB FOA are indicated by black, red and blue solid lines, respectively); (c) scattering based fixed attenuators principles (the curves for 20 dB FOA, 15 dB FOA and 5 dB FOA are indicated by black, red and blue solid lines, respectively).
@@
Due to variations in efficiency of coupling to the single-mode fiber, losses due to absorption by the filters and technical restrictions imposed by the spectrometer, the measured spectrum of the broadband light source (BLS) was narrowed to the 400-800 nm range. Insertion losses of the optical element under test was calculated as follows:
@@
III Experimental results
III.1 Optical attenuators
In this section, we present the results for two variable optical attenuators (VOAs) that differ in the mode of operation along with six different fixed optical attenuators (FOAs).
In QKD systems optical attenuators (OA) are typically employed to reduce the intensity of laser pulses to the single photon level. Additionally, OAs are of importance in reducing the probability of wavelength-dependent attacks successful realization. Consequently, there is a requirement to avoid pronounced variations of spectral insertion loss properties of OAs across a broad wavelengths range.

### v3

@@ Abstract
In this paper we investigate spectral vulnerabilities in quantum key distribution systems arising from the use of shorter-wavelength radiation in the 400–800 nm range, with particular focus on the induced photorefraction attack (IPA). Crucial elements influenced by IPA include various types of modulators, both phase and intensity modulators. In the following paper, we consider different scenarios and their implications. Through combined theoretical and experimental analysis, we demonstrate that optical components commonly used as countermeasures in the telecom band (1000–2100 nm) exhibit significantly reduced effectiveness at shorter wavelengths. The efficiency of IPA is shown to increase as the wavelength decreases, posing a substantial threat to phase-modulation-based QKD protocols. We analyze the impact of IPA across different QKD architectures and assess the feasibility of potential countermeasures under realistic implementation scenarios. Our results highlight the necessity of broadband security evaluations and wavelength-aware component design in future QKD systems.
@@ I Introduction
The security of quantum key distribution (QKD) systems hinges upon the fundamental laws of physics. Over the past decades, significant progress has been made in theoretical security proofs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. However, many of these proofs rest on the assumption of perfect and trusted devices. Such trust assumption imposes a rather strong security requirement, which is generally difficult to implement in practice. In real situations, imperfect devices in QKD systems would introduce vulnerabilities that can be exploited by eavesdroppers to extract additional information about the transmitted bits during their attacks, as it is considered in [11, 12, 13, 14, 15, 16]. Device-independent (DI) [17, 18, 19, 20, 21, 22] and semi-device-independent [23, 24, 25, 26] protocols have emerged as a prominent area of research that provides a way to prevent side-channel attacks [27]. On the opposite side, another rapidly developing field which is often referred to as quantum hacking has received a considerable amount of attention [28]. The field of quantum hacking focuses on identifying and addressing vulnerabilities in the technical implementation of QKD systems. The significance of researching hardware loopholes in QKD systems arises from their transition from laboratory prototypes to commercial deployment. As QKD technologies advance toward practical implementation, identifying and mitigating potential hardware vulnerabilities becomes crucial to ensuring the security and reliability of real-world quantum communication networks [29, 30, 31, 32].
@@
Emerging approaches in quantum hacking pose novel challenges for security analysis of QKD systems. Numerous vulnerabilities have been identified in prior works [33, 34, 35], underscoring the importance of enhancing security measures [36, 37, 38, 39]. Eavesdropper strategies in quantum hacking can broadly be classified under two basic types, active [40, 41, 42] and passive [43, 44], as well as combinations thereof. Active strategies involve the manipulation of optical parameters within the hardware, whereas passive strategies solely facilitate the acquisition of additional information.
@@
A quintessential illustration of a passive attack is the Trojan-horse attack (THA), which is also known as the large pulse attack [45, 44]. In this scenario, the eavesdropper (Eve) employs scanning pulses introduced into the systems of legitimate users (Alice and Bob) via the quantum channel, allowing for the extraction of distributed sequences from the reflected portions of the eavesdropper pulses back into the quantum channel. The key factors that determine the efficiency of this attack involve the transmittance of elements in two directions, the distinguishability of states along with countermeasures utilized in the system. An important point is that all these factors require the eavesdropper to define the optimal spectral range of wavelengths employed to perform attack.
@@
The well-known detector blinding attack (DBA) [46, 47, 48] exemplifies an active attack. The idea underlying the attack is that, under the action of high-power laser radiation (as compared to typical intensities used in QKD), the operating mode of single photon detectors (SPDs) based on avalanche photodiodes can be changed from the Geiger mode into the linear operating mode, so that the eavesdropper can manipulate the clicks registered by the SPDs. The efficiency of this DBA based quantum hacking approach crucially depends on the spectral sensitivity of the avalanche photodiode, the transmittance of optical elements connecting the quantum channel with the single photon avalanche diode (SPAD), and the countermeasures implemented into the QKD system. So, it turned out that, similar to the Trojan-horse attack, the spectral range is one of the factors that govern the success of the DBA attack.
@@
Previous studies [49, 50, 51, 52, 53] motivated by the THA, have mainly been focused on transmittance of fiber optical elements utilized in QKD systems at wavelengths ranged between 1000 nm and 2100 nm. For the 600-1000 nm range, the experimental results were reported in Refs. [52, 54].
@@
Note that, according to Ref. [60], successful implementation of the IPA requires low intensities starting from 3 nW. A comparison between these intensities and the maximal intensity permissible for insertion into a fiber without causing damage, which is typically around 9 W (this value is applicable to quartz single mode fibers without power damage level doping commonly used in fiber communications and may vary among manufacturers) leads to the conclusion that attenuation needed for passive defence against IPA for single pass through optical elements can roughly be estimated at about 100 dB. Thus, in order to assess the feasibility of the attack, it is necessary to measure and analyze insertion losses of individual optical elements on the side of Alice. To this end, we study the insertion losses of different optical elements in the 400–800 nm spectral range, demonstrate the existence of potential vulnerabilities and suggest potential countermeasures.
@@
The paper is organized as follows. Security issues arose due to considered effect are discussed in Sec. II. In Sec. III we describe our experimental technique employed to measure spectra of insertion losses of optical elements at wavelengths ranged from 400 nm to 800 nm. The experimental data for insertion losses measured for a variety of optical attenuators, wave-division multiplexing (WDM) and isolating components are presented in Sec. IV. Sec. V is dedicated to countermeasures discussion. Finally, Sec. VI concludes the article.
II Security issues
In considerations of side-channel attacks aimed at extracting quantum information, the visible spectral range is often disregarded on account of its poor confinement in single-mode fibers. Nevertheless, this range plays a significant role in the manipulation of optical components. Techniques such as laser damage and induced photorefraction exploit visible-wavelength radiation to impair or alter the functionality of critical devices without directly accessing the quantum channel.
In this work, we focus on the induced-photorefraction attack (IPA) as a representative example of such vulnerabilities. We assess its potential impact across a range of QKD protocols and system architectures described in the literature. Particular focus is placed on phase and intensity modulators, whose disruption poses a direct threat to state preparation fidelity. This section highlights these vulnerabilities and outline their broader consequences for QKD system security.
Change of phase modulator’s (PM’s) modulation index may be considered separately, depending on implemented optical devices. A PM can be a part of the following commonly used in the field of QKD devices:
  1. 1.
    PMs introducing a constant phase shift within Mach–Zehnder interferometer (MZI) schemes;
  2. 2.
    PMs with an oscillating phase shift, as utilized in subcarrier wave (SCW) QKD [13, 33, 61, 21];
  3. 3.
    PMs integrated into variable optical attenuators (VOAs);
  4. 4.
    PMs employed as key elements in amplitude modulators (AMs) or intensity modulators (IMs).
In addition to the above classification, QKD protocol and utilized quantum states play essential role in security estimations. Accordingly, we distinguish the following categories of QKD schemes:
  1. 1.
    Protocols utilizing linearly dependent states (such as single-photon BB84 [62] or single-photon fraction in decoy-state protocols [12, 63, 64]), where possible underestimation of key parameters or phase-remapping attack [65] should be taken into consideration. The security problem for such protocols has already been studied in a number of papers[60, 59, 28];
  2. 2.
    Protocols with linearly independent set of states (coherent states with phase-coding [66, 67, 68, 69] or SCW QKD have not been considered so far in the light of IPA auxiliary attack. In what follows, we concentrate on these kinds of protocols.
II.1 Linearly independent set of states
Now the second group of the protocols listed above will be our primary concern. We begin with the protocols utilizing coherent states with phase-coding[66, 67, 68, 69].
Roughly speaking, when a nonlinear material such as lithium niobate is illuminated with the pump light, the photorefractive effect gives rise to the light-induced change of the refractive index arising through the linear electro-optic (Pockels) effect from the space-charge electric field caused by inhomogeneous redistribution of photogenerated charges (see, e.g., [28]; and references therein for more details). The photorefractively changed refractive index leads to the photoinduced shift of the signal phase that, in turn, affects electro-optic response and the working states of lithium niobate-based devices widely employed in QKD systems.
In particular, owing to this phase shift, the voltage modulation curves of MZI-based VOA and IM appear to be displaced along the voltage axis introducing deviations of the DC-bias voltage. This effect has been previously employed in the context of IPAs against the measurement-device-independent (MDI) QKD [60] the decoy-state BB84 QKD [28] and the Gaussian-modulated coherent-state continuous-variable QKD in [70].
In our case, however, the assumption that the only photorefraction-induced effect is the phase shift identical for all states does not provide any advantage to an eavesdropper. The reason is that all the states are equally distributed in the phase plane and such phase shift does not affect the phase difference between the states.
From the other hand, one of the experimentally measured results reported in [59] is the photorefractively induced change of the half-wave voltage in PMs that can be described as the photoinduced change of the modulation index. According to Ref. [59], this effect is similar to the phase-remapping attack [65], causing the phase difference between coded states to change.
In this case, phase-remapping by a factor of $x$ of is applicable:
$\{|\alpha_{0}\rangle,\ |\alpha_{0}e^{\frac{i\pi}{N}}\rangle,\ |\alpha_{0}e^{\frac{i2\pi}{N}}\rangle,\ ...\ ,\ |\alpha_{0}e^{\frac{i(2N-1)\pi}{N}}\rangle\}\rightarrow$
$\rightarrow\{|\alpha\rangle,\ |\alpha e^{\frac{ix\pi}{N}}\rangle,\ |\alpha e^{\frac{i2x\pi}{N}}\rangle,\ ...\ ,\ |\alpha e^{\frac{i(2N-1)x\pi}{N}}\rangle\},$ (1)
where $\alpha_{0}$ is the complex amplitude of the coherent state, $\alpha$ is the complex amplitude after IPA with $|\alpha|\leq|\alpha_{0}|$, $2N$ is the total number of states in the set. After the IPA, the eavesdropper can apply additional commonly used attack, e.g. unambiguous state discrimination (USD) attack.
This attack is characterized by the USD probability, $P_{U}(\alpha,x)$, giving the optimal probability for USD of a set of states. When the states are equiprobable, this probability can be evaluated by computing the minimal eigenvalue of the Gram matrix, $G$ [71, 72, 73, 13]. For the unperturbed coherent states, which are symmetrically arranged, the eigenvalues can be obtained in the closed form using the discrete Fourier transform technique (see, e.g. [73]). After the IPA, the Gram matrix, which is a positive semidefinite Toeplitz matrix, is no longer represented by a circulant matrix and computing the eigenvalues requires numerical analysis.
However, a few analytical observations can be made. In particular, we have found that the ratio of the probabilities $f(x)=P_{U}(\alpha,x)/P_{U}(\alpha,1)$ does not exceed the unity, $0\leq f(x)\leq 1$. In addition, for $N>1$, it weakly depends on $|\alpha|$. For the special case with just two states as in B92 and $N=1$, we have
$P_{U}(\alpha,x)=1-e^{-|\alpha|^{2}(1-\cos(\pi x))}$ (2)
and
$f(x)=\frac{1-e^{-|\alpha|^{2}(1-\cos(\pi x))}}{1-e^{-2|\alpha|^{2}}}.$ (3)
The results for the USD probability ratio computed as a function of $x$ at $N=2$, $N=3$ and $N=4$ are shown in Fig. 1. Also note, that
$P_{U}(\alpha,1)\approx\frac{2N(|\alpha|^{2})^{2N-1}}{(2N-1)!}.$ (4)
The overlaps of the states after the IPA only increase, so it cannot provide any significant advantage to an eavesdropper.
Figure 1: USD probability ratio $f(x)=P_{U}(\alpha,x)/P_{U}(\alpha,1)$ computed as a function of the phase-remapping parameter $x$ (see Eq. (1)) at (a) $N=2$, (b) $N=3$, (c) $N=4$. The curves show that phase-remapping with non-integer $x$ leads to a significant reduction in unambiguous state discrimination probability.
II.2 Intensity variation in SCW QKD
As for SCW QKD setup, IPA here plays the same role as attacking an IM or a VOA based on a PM in protocols with coherent states and phase-coding, since total power of sidebands is a function of modulation index $m$ , more precisely,
$|\alpha|^{2}=|\alpha_{0}|^{2}(1-J_{0}^{2}(m))\approx\frac{|\alpha_{0}|^{2}m^{2}}{2},$ (5)
where approximation is valid for small values of $m<1$ and $J_{0}(m)$ is Bessel function of the first kind), and $|\alpha_{0}|^{2}$ is the total power. And the modulation index can be increased by utilizing the IPA with a recovery, as described in [59], where legitimate parties recalibrate their setups after IPA. In turn, an increase of the modulation index significantly enhances the information accessible to an eavesdropper, as it can be seen in Fig. 2. This quantity is known as Holevo bound, and it can be formally evaluated as follows [61]:
$\chi(A:E)=h\Big{(}\frac{1-e^{-|\alpha_{0}|^{2}(1-J_{0}^{2}(2m))}}{2}\Big{)}.$ (6)
Figure 2: Variation of Holevo bound $\chi(A:E)$ as in Eq. (6) depending on the (multiplicative) factor $\Delta m$ of modulation index $m$, denoted as solid black line, compare to the initial value of the Holevo bound with $\Delta m=1$, denoted as dashed gray line; calculations were performed for the following values: $|\alpha_{0}|^{2}=1$, $m=0.434$ that determines the ratio between the power of the sidebands to the power of the carrier to be equal $1:10$
II.3 Intensity variation in decoy-state
A similar concern emerges in the implementation of decoy-state QKD protocols [12], where precise parameter estimation—particularly of the total gain $Q_{\mu}$, the photon-number-dependent yield $Y_{n}$, single-photon error fraction $E_{1}$ and total error $E_{\mu}$ —is vital for security proofs. These parameters depend on the mean photon number $\mu$, for which tight bounds are required. However, most security frameworks assume that $\mu$ remains stable and immune to external manipulations. However, under the influence of induced photorefraction, this assumption may no longer hold. The external manipulation of optical intensity by an adversary can lead to an artificial increase in $\mu$, resulting in erroneous parameter estimation and potential security compromise, as discussed in [59, 60].
Additionally, such intensity modifications open the door to phase-remapping attacks [65]. Despite proposed approaches were reported to be suboptimal, they are feasible with current technological level.
These considerations highlight the protocol- and implementation-specific nature of IPA vulnerabilities. As a result, a comprehensive security analysis must include accurate optical characterizations across both infrared and visible spectral regimes.
III Experimental approach
While the previous section addressed the general security implications of IPA, it is essential to recognize that such effects manifest only under well-defined optical conditions. A realistic assessment of IPA threats in QKD systems requires quantifying the probe light power that can reach susceptible components, particularly LiNbO3-based modulators. In general, assuming that the maximal power of radiation that can propagate in silica optical fibers without damage is about 40 dBm [74, 75, 56], the above power of probing radiation can be evaluated using the following expression:
$P_{\mathrm{dBm}}(\lambda)=40-\sum_{i=1}^{n}(\alpha_{\mathrm{n}}(\lambda)),$ (7)
where $P_{\mathrm{dBm}}(\lambda)$ is the estimated power of Eve’s source which will reach the modulator under the attack, an ($\lambda$) are insertion losses of the nth fiber-optical element (the elements are ordered along the path of Eve’s radiation).
Clearly, using this formula requires a systematic characterization of the optical components’ response in the visible spectral range. This section presents an experimental framework for evaluating these characteristics and examines the corresponding implications for system design and countermeasure implementation.
@@ III Experimental approach
Figure 3 shows the optical scheme of our experimental setup used to measure the insertion losses of optical elements in the visible range. In this scheme, a broadband plasma light source (XWS65, ISTEQ) indicated as the broadband light source (BLS) is utilized to produce light with wavelengths ranged from 250 nm up to 2500 nm. Referring to Fig. 3, the radiation from this source is directed into a single-mode fiber ITU-T G.652 (OF) passing through a set of lenses (L), filters (F) and the objective (O) which is positioned on a three-axis stage to meet coupling and spectrum requirements. After the fiber, the focused light propagates through either a fiber or a fiber with the optical elements under test and is measured by the spectrometer (USB4000-UV-VIS-ES, Ocean Optics) with the 1.5 nm resolution.
@@
Due to variations in efficiency of coupling to the single-mode fiber, losses due to absorption by the filters and technical restrictions imposed by the spectrometer, the measured spectrum of BLS was narrowed to the 400-800 nm range. Insertion losses of the optical element under test were calculated as follows:
@@
The optical scheme behind our experimental technique enables us to perform measurements of the insertion losses with approximately $50$ dB dynamic range. It is simple to use and low-cost when compared with analogous schemes discussed, e.g., in Ref. [35]. An important point is that, by using suitably modified methods presented in Ref. [51], the proposed scheme can be improved so as to extend both the wavelength range and the dynamic range of measurements. As far as certification of QKD systems is concerned, this is a feasible and easy-to-use scheme that can be employed for characterization of optical elements’ insertion losses.
IV Experimental results
IV.1 Optical attenuators
In this section, we present the results for two variable optical attenuators (VOAs) that differ in the mode of operation, along with six different fixed optical attenuators (FOAs).
In QKD systems, optical attenuators (OA) are typically employed to reduce the intensity of laser pulses to the single photon level. Additionally, OAs are of importance in reducing the probability of wavelength-dependent attacks successful realization. Consequently, there is a requirement to avoid pronounced variations of spectral insertion loss properties of OAs across a broad wavelengths range.
@@ IV.1 Optical attenuators
Figure 4: Insertion losses measured in the 400-800 nm range for (a) electro-mechanical and (b) electro-optical variable optical attenuators (blue and red solid lines indicate the zero-voltage curve and the case where the applied voltage is 5 V, respectively); (c) absorption based fixed attenuators (the curves for the first 20 dB FOA, second 20 dB FOA and 10 dB FOA are indicated by black, red, and blue solid lines, respectively); (d) scattering based fixed attenuators principles (the curves for 20 dB FOA, 15 dB FOA and 5 dB FOA are indicated by black, red, and blue solid lines, respectively).
In the zero-voltage case, it turned out that the measured insertion losses are higher than the values reported in the datasheet. For EM VOA and EO VOA, the minimal losses 10.7 dB and 21.8 dB are reached at the wavelengths 761 nm and 557 nm, respectively.
Referring to Figs. 4(a) and (b), when the applied voltage is 5 V (blue lines), the losses are shown to vary in the range between 13.7 dB and 30 dB for EM VOA, whereas, for EO VOA, they are ranged from 23.7 dB to 50 dB. Note that these maximum values of losses are approximate, as they correspond to the noise level of our experimental setup and depend on the number of neutral filters.
Insertion losses measured in relation to the wavelength for six FOAs are shown in Figs. 4(c) and (d). For three absorption based FOAs (see Figs. 4(c)), according to the values reported by the manufacturers, the insertion losses of one FOA is 10 dB, whereas the losses for the other two FOAs are the same and equal to 20 dB. For the scattering based FOAs (see Figs. 4(d)), there are three different reported values of losses at the operating wavelength ($\lambda=1550$ nm): 5 dB, 15 dB and 20 dB. These values are in good agreement with the results of our measurements performed at $\lambda=1550$ nm.
The insertion losses of the absorption based FOAs measured in the 400-800 nm range are shown in Fig. 4(c). It can be seen that, for each FOA, there are two transmittance windows where the insertion losses are noticeably lower than the value at the operating wavelength indicated in the legend. The window ranged between 400 nm and 430 nm is located in the vicinity of local minima for the two 20 dB attenuators with $\alpha_{\mathrm{dB}}\approx 8.1$ and $\alpha_{\mathrm{dB}}\approx 8.6$ at the wavelength $405$ nm (for the 10 dB attenuator, the corresponding value is $\alpha_{\mathrm{dB}}\approx 2.7$) Similarly, the other window in the 750-800 nm wavelengths range is close to the local minima at $\lambda\approx 780$ nm with the insertion losses estimated at $\alpha_{\mathrm{dB}}\approx 3.4$ and $\alpha_{\mathrm{dB}}\approx 4.3$ and $\alpha_{\mathrm{dB}}\approx 0.9$.
Figure 4(d) shows the wavelength dependencies of the insertion losses measured for the scattering based FOAs. It is seen that, for the 20 dB attenuator, the insertion losses insignificantly change with wavelength, varying from 27 dB to 25 dB. A similar conclusion applies to the 15 dB (5 dB) attenuator, where variations are between 15.6 dB (11.2 dB) and 13 dB (10.8 dB).
IV.2 Wavelength-division multiplexing components